OpenStudy (anonymous):

Given the following perfect square trinomial, fill in the missing term. x2 − 16x + ____

OpenStudy (cwrw238):

square half of the coefficient of x

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

since (x+y)^2=x^2+y^2+2yx, 2yx=-16x, and y =-8, and ___=(-8)^2 =64

OpenStudy (cwrw238):

the coefficient of x is -16 so its -8 ^2 = 64

OpenStudy (anonymous):

ooooooooh okay, thank you!

OpenStudy (cwrw238):

yw